OpenStudy (anonymous):

A. c = 5 B. c = 6 C. c = 150 D. c = 180

hero (hero):

You posted the answer choices. Where's the question?

OpenStudy (anonymous):

c • 5 = 30

OpenStudy (anonymous):

Use transformations to solve the equation.

OpenStudy (anonymous):

B. c=6 because to find c you divide 30 by 5 30/5 = 6

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

–9b = –45

OpenStudy (anonymous):

Unit 3 | Lesson 8 Lesson Assessment: Equations Involving Multiplication and Division Quiz back 2. Use transformations to solve the equation. –9b = –45 A. b = –9 B. b = –5 C. b = 5 D. b = 9 Question Resources Advance Preparation Hints Question 2 of 5Materials Objectives Only text allowed

hero (hero):

In order to solve this problem, you will need to know how to: 1. isolate variables 2. inverse operations Which of these are you unfamiliar with?

OpenStudy (anonymous):

isolate variables

hero (hero):

When you isolate a variable in an equation, you perform operations such as addition, subtraction, multiplication or division, such that only the variable itself remains on one side of the equation while the rest of the terms and expressions are presented on the other side of the equation. For example, if we isolated x in the equation 2x + 4y = 10, it would become: \[x = \frac{10 - 4y}{2}\] In this form x has been isolated using the inverse operations subtraction and division.